40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1596     Entry time: Sun Oct 25 22:15:35 2015
Author: Antonio 
Type: Summary 
Category: EOM 
Subject: AEOM installed in the South path. Noise issue. 

Summary

The AEOM has been installed in the South path replacing the EOM 21MHz used for the PMC. There is a high noise that I clearly see at the photodiode in transmission. 

 

When I have placed the AEOM in the path I have decided to take the alignment of the previous EOM as reference. Not ideal because the reference should be the incoming beam. The beam is not parallel to the table and it was decided to be as less as possible invasive. The mode matching and the alignment gave at that time 20% of visibility (at each polarization). After the installation parameters where unchanged. Later I have improved the alignment bringing the visibility at 30% for both the polarizations. After that, when everything was in place I have easily locked the cavity but the power in transmition was showing a very high noise. I have spent all the day trying to twick the alignment because and servo loop gain, but we need to solve this before going further. My back does not allow me to proceed for today.

 

NOTE:

I have also noted that the South Laser which is labeed 2W laser has the lambda/4 and the lambda/2 rotated in a way that at the output of FI we had few mm. I am not sure if damping the power at the FI is a good thing.

ELOG V3.1.3-